OpenStudy (anonymous):

Simulating a problem, please help! I am asked to simulate the probability of statements like "collect all 8 dinosaurs" or "one in each specially marked cereal box." I spun a wheel 3 times, and my answers were these for the 1st, 2nd, and 3rd spin: Tyrannosaurus rex: 2, 1, 4 Velociraptor: 3, 1, 1 Apatosaurus: 1, 1, 5 Stegosaurus 3, 2, 1 Pterodactyl: 2, 2, 2 Allosaurus: 3, 3, 1 Gigantosaurus: 1, 2, 1 Triceratops: 2, 3, 2 I have no idea how to solve this...

OpenStudy (anonymous):

The first question I have to answer is: What was the theoretical probability of collecting a specific dinosaur, such as a Velociraptor, in a single purchase?

OpenStudy (dan815):

what do those numbers represent when u spin

OpenStudy (anonymous):

The spinner is to determine how many boxes of cereal I might need to purchase to collect all eight dinosaurs.

OpenStudy (dan815):

ok

OpenStudy (dan815):

Tyrannosaurus rex: 2, 1, 4 -->calculate the avg of the boxes u need to purchase Velociraptor: 3, 1, 1 Apatosaurus: 1, 1, 5 Stegosaurus 3, 2, 1 Pterodactyl: 2, 2, 2 Allosaurus: 3, 3, 1 Gigantosaurus: 1, 2, 1 Triceratops: 2, 3, 2

OpenStudy (anonymous):

Do I add, divide, or multiply?

OpenStudy (dan815):

tell me this say you are 12 years old and your brother is 10 years old what is the average age between you 2

OpenStudy (anonymous):

2 years

OpenStudy (dan815):

oh, u dont know what average means?

OpenStudy (dan815):

lets say you write 2 tests you get 50% on one and you get 100% on the other what is your average right now?

OpenStudy (anonymous):

Oh, ok, my average then is 75%, correct? Sorry, I was a bit confused

OpenStudy (dan815):

yes

OpenStudy (dan815):

so ur telling me each time you spin that tells you how many boxes you needed to get that dionsaur

OpenStudy (dan815):

so your needed 2 boxes on the first try to get that dinosaur and 1 boxes the next time to get that dinosaur and 4 boxes the 3rd time to get that dinosaur, so on average how often are you getting that dinosaur

OpenStudy (anonymous):

Ah, 2.33?

OpenStudy (dan815):

yeah ok

OpenStudy (dan815):

that is how the spinnner is right

OpenStudy (dan815):

can you tell me your question as it is

OpenStudy (dan815):

"The spinner is to determine how many boxes of cereal I might need to purchase to collect all eight dinosaurs." are you sure they* dont mean those 3 spins together determine how many boxes you need to get that dinosaur?

OpenStudy (anonymous):

"For this exercise, the probabilities refer to the probability of getting any one dinosaur when making any one purchase. You do not need to determine the probability of the compound event of getting all eight of the dinosaurs."

OpenStudy (dan815):

ok

OpenStudy (dan815):

i think u those 3 spins are saying how many boxes in total u need to get that 1 dinosaur so just add them

OpenStudy (dan815):

i dont think theyll make it so u have to calculate avg for everyone.. thats a little be too tedious and the probability values dont make sense

OpenStudy (dan815):

Tyrannosaurus rex: 2, 1, 4 =7 Velociraptor: 3, 1, 1=5 Apatosaurus: 1, 1, 5=7 Stegosaurus 3, 2, 1=6 Pterodactyl: 2, 2, 2=6 Allosaurus: 3, 3, 1=7 Gigantosaurus: 1, 2, 1=4 Triceratops: 2, 3, 2=7

OpenStudy (dan815):

Tyrannosaurus rex: 2, 1, 4 =7 ; Prob = 1/7 Velociraptor: 3, 1, 1=5; prob = 1/5 Apatosaurus: 1, 1, 5=7; p=1/7 Stegosaurus 3, 2, 1=6; p= 1/6 Pterodactyl: 2, 2, 2=6 p = 1/6 Allosaurus: 3, 3, 1=7 p = 1/7 Gigantosaurus: 1, 2, 1=4 p = 1/4 Triceratops: 2, 3, 2=7 p= 1/7

OpenStudy (anonymous):

Ahhh, now I think get it, thanks a lot!